OpenStudy (anonymous):

A cannery processed 1965 pounds of strawberries in 6.5 hours. The cannery processed 3365 pounds in 10.5 hours. A. Write a linear equation to model the weight of strawberries S processed in T hours. B. How many pounds of strawberries can be processed in 12 hours?

OpenStudy (anonymous):

Because only parameters are given, let us assume the relation as linear Let S(T)= AT+B where S(T) is the weight of strawberries processed in time T 1,445 pounds of strawberries in 4.5 hours →1,445=A× 4.5 +B 2,320 pounds in 7 hours→ 2320=A× 7+B Solving we get A= 350 and B= −130 Hence S(T)= 350T−130 Strawberries can be processed in 11 hours S(11)= 350×11−130=3720 pounds

OpenStudy (anonymous):

Thank you Gerardie. Helped a lot.

OpenStudy (anonymous):

No problem bro